What is the greatest number of rectangles with integer side lengths and perimeter 10 that can be cut from a piece of paper with width 24 and length 60?
A)144
B)180
C)240
D)360
E)480

Answer by htmentor(789)

The maximum number of rectangles will be obtained if the area of each cut rectangle is a minimum
The perimeter of each cut rectangle = P = 2(l+w) = 20 -> l+w = 5 -> w = 5-l
The area of each cut rectangle = A = l*w = l*(5-l)
The possible values for l are 1,2,3 and 4, giving areas of 4, 6, 6 and 4
So l=4 gives the minimum area (letting the length be the longer dimension)
Take the ratio of the areas to give the number of cut rectangles:
24*60/4 = 6*60 = 360
Ans: D